# and cross estimates

### v1

@@ 1.4. Examples
[…]
@@
[…]
@@
[…]
@@
[…]
@@ 2.1. Mean-field relaxation
For completeness, we include a short proof of item (i) as a consequence of (ii) for small $\kappa$. Let $\mu,\mu^{\prime}\in C(\mathbb{R}_{+};\mathcal{P}(\mathbb{X})\cap\operatorname{L}^{2}(\omega))$ satisfy the mean-field equation (1.8) with initial data $\mu_{\circ},\mu_{\circ}^{\prime}\in\mathcal{P}(\mathbb{X})\cap\operatorname{L}^{2}(\omega)$. Treating interactions perturbatively, Duhamel’s formula for (1.8) yields
@@
For completeness, we emphasize that the assumptions in [14] are indeed satisfied for the operator $L_{0}$ on $L^{2}(\omega)$ with $A$ satisfying (1.3), so we can apply their hypocoercivity results and deduce (2.4) above. Assumption (H1) in [14] is simply the Gaussian Poincaré inequality, and assumption (H3) is trivial for kinetic equations. Next, as $A$ is uniformly convex, we recall that the probability measure $\propto e^{-A}$ on $\mathbb{R}^{d}$ satisfies a Poincaré inequality, see e.g. [32, Section A.19]. By [13, Lemma 3], this ensures the validity of Assumption (H2) in [14]. Finally, given $w\in C^{\infty}_{c}(\mathbb{R}^{d})$, for the solution $u$ of
@@
We recall the following useful observation, see e.g. [17, Lemma II.3.10]: if $(T_{t})_{t\geq 0}$ is a semigroup on a Banach space $(B,\|\cdot\|_{B})$ and if there exist $P>0$ and $\omega\in\mathbb{R}$ such that we have $\|T_{t}h\|_{B}\leq Pe^{\omega t}\|h\|_{B}$ for all $t\geq 0$ and $h\in B$, then there exists a modified norm $\operatorname{|\hskip-1.13809pt|\hskip-1.13809pt|}\cdot\operatorname{|\hskip-1.13809pt|\hskip-1.13809pt|}_{B}$ on $B$ that is Lipschitz-equivalent to $\|\cdot\|_{B}$ such that $\operatorname{|\hskip-1.13809pt|\hskip-1.13809pt|}T_{t}h\operatorname{|\hskip-1.13809pt|\hskip-1.13809pt|}_{B}\leq e^{\omega t}\operatorname{|\hskip-1.13809pt|\hskip-1.13809pt|}h\operatorname{|\hskip-1.13809pt|\hskip-1.13809pt|}_{B}$, where now the multiplicative constant is equal to 111. Indeed, it suffices to consider the norm
@@
where we used that the orthogonal projection $\pi^{\bot}$ commutes with the semigroup. For $m\geq 1$, we also define $\operatorname{|\hskip-1.13809pt|\hskip-1.13809pt|}\cdot\operatorname{|\hskip-1.13809pt|\hskip-1.13809pt|}_{L^{2}(\omega^{\otimes m})}$ as the associated injective cross norm on $\operatorname{L}^{2}(\omega^{\otimes m})$,
@@ 3. For $P\subset\llbracket m\rrbracket$ and $k,\ell\in P$, we define the operators
    $S_{k,\ell}h^{P}$ ::=:= $-K(x_{k},x_{\ell})\cdot\nabla_{v_{k}}h^{P},$
    ${}\displaystyle H_{k}h^{P\cup\{*\}}$ ::=:= $-\int_{\mathbb{X}}K(x_{k},x_{*})\cdot\nabla_{v_{k}}h^{P\cup\{*\}}\,\mathrm{d}z_{*}.$
@@ 2.3. Gibbs relaxation
[…]
@@
Now appealing to (2.7) (with constant 111) and to (2.2) (with $L_{0}$ replaced by $L_{0}^{*}$ and using the equivalence of norms (2.7)), the claim (2.12) follows.
@@ 2.4. Correlation functions
where the symbol ‘${\operatorname{sym}}$’ stands for the symmetrization of coordinates, and we can write more generally, for all $1\leq m\leq N$,
@@ 2.5. BBGKY hierarchy for correlations
Starting from the BBGKY hierarchy (2.9) for marginals of $F^{N}$, and recalling how marginals can be expanded in terms of correlations (2.15) and vice versa (2.17), we can derive a corresponding BBGKY hierarchy of equations for correlations. We refer for instance to [23, Section 4] for the derivation: straightforward but lengthy algebraic manipulations are needed to collect the different factors in this hierarchy.
@@
Given $K\in L^{\infty}(\mathbb{R}^{2d})^{d}$, consider a weak solution $F^{N}\in C(\mathbb{R}_{+};\mathcal{P}\cap\operatorname{L}^{1}(\mathbb{X}^{N}))$ of the Liouville equation (1.4). The associated correlation functions $\{G^{N,m}\}_{1\leq m\leq N}$ satisfy the following hierarchy of equations: for all $1\leq m\leq N$,
@@ 2.6. Uniform-in-time propagation of chaos and correlation estimates
${}.(\partial_{t}-L_{0})(F^{N,1}-\mu)\,=\,\kappa H_{1}\Big{(}\frac{N-1}{N}F^{N,\{1\}}F^{N,\{*\}}-\mu(z_{1})\mu(z_{*})\Big{)}+\kappa\frac{N-1}{N}H_{1}G^{N,\{1\}\cup\{*\}}.$
@@ 3.1. Relaxation of correlations
[…]
@@ 4.1. Mean-field relaxation
As $A$ is uniformly convex, cf. (1.3), we recall that the probability measure $\propto e^{-A}$ satisfies a Poincaré inequality, see e.g. [32, Section A.19]: in terms of the weight $\omega=e^{A}$, this reads as follows, for all $h\in C^{\infty}_{c}(\mathbb{R}^{d})$,
@@ 4.2. BBGKY hierarchy
$S_{k,\ell}h^{P}$ ::=:= $-{\operatorname{div}}_{x_{k}}(K(x_{k},x_{\ell})h^{P}),$
${}\displaystyle H_{k}h^{P\cup\{\ast\}}$ ::=:= $-{\operatorname{div}}_{x_{k}}\Big{(}\int_{\mathbb{R}^{d}}K(x_{k},x_{*})h^{P\cup\{\ast\}}\,\mathrm{d}x_{*}\Big{)}.$ (4.3)
[…]
@@ 5.1. Ergodic estimates for mean-field operators
$L_{M}h\,=\,\left\{\begin{array}[]{lll}\triangle h&:&\text{in Case\leavevmode\nobreak\ 1},\\ \triangle h+\kappa\triangle W\ast h&:&\text{in Case\leavevmode\nobreak\ 2}.\end{array}\right.$
@@ 5.3. Relaxation of correlations
[…]

### v2

@@ 1.5. Examples
We briefly describe some applications of our results. In item (b), we give an example with merely bounded interaction forces, for which even our uniform-in-$N$ Gibbs relaxation result in Proposition 1.1(ii) is new.
@@
    for some fixed $\sigma>0$. In this setting, both for the underdamped and for the overdamped Langevin dynamics, all our results apply: for weak interactions, that is, for $\kappa$ small enough, Proposition 1.1 yields uniform-in-time correlation estimates and uniform-in-$N$ Gibbs relaxation, and Theorem 1.2 further yields new estimates for the cross mean-field/relaxation error.
@@
    for instance with the periodized Gaussian kernel $\Gamma_{\sigma}(x)=(2\pi\sigma^{2})^{-d/2}\sum_{n\in\mathbb{Z}^{d}}\exp(-\frac{1}{2\sigma^{2}}|x+n|^{2})$, for some fixed $\sigma>0$. For $\kappa$ small enough, the same results hold in this setting, and we recall that Corollary 1.3 then further yields an improved convergence rate for the one-particle density (see Corollary 1.5 in the overdamped setting). In addition, it is easily checked that $K$ satisfies the assumptions of Case 2 in Section 1.4: hence, in the overdamped setting, without any smallness requirement on $\kappa>0$, Theorem 1.6 yields a uniform-in-$N$ Gibbs relaxation result and cross error estimates up to $O(N^{-\infty})$ in negative Sobolev norms.
@@
    for some fixed radius $r>0$. In this setting, for $\kappa$ small enough, uniform-in-time propagation of chaos and correlation estimates follow from [34]. Corollary 1.5 further provides an improved convergence rate for the one-particle density. Next, since $K$ is not Lipschitz, we obtain the first uniform-in-$N$ Gibbs relaxation result, cf. Proposition 1.1(ii). Finally, Theorem 1.4 provides new estimates on the cross mean-field/relaxation error.
@@ 2.1. Mean-field relaxation
For completeness, we include a short proof of item (i) as a consequence of (ii) for small $\kappa$. Let $\mu,\mu^{\prime}\in C(\mathbb{R}_{+};\mathcal{P}(\mathbb{X})\cap{L}^{2}(\omega))$ satisfy the mean-field equation (1.8) with initial data $\mu_{\circ},\mu_{\circ}^{\prime}\in\mathcal{P}(\mathbb{X})\cap{L}^{2}(\omega)$. Treating interactions perturbatively, Duhamel’s formula for (1.8) yields
@@
For completeness, we emphasize that the assumptions in [14] are indeed satisfied for the operator $L_{0}$ on $L^{2}(\omega)$ with $A$ satisfying (1.3), so we can apply their hypocoercivity results and deduce (2.4) above. Assumption (H1) in [14] is simply the Gaussian Poincaré inequality, and assumption (H3) is trivial for kinetic equations. Next, as $A$ is uniformly convex, we recall that the probability measure $\propto e^{-A}$ on $\mathbb{R}^{d}$ satisfies a Poincaré inequality, see e.g. [33, Section A.19]. By [13, Lemma 3], this ensures the validity of Assumption (H2) in [14]. Finally, given $w\in C^{\infty}_{c}(\mathbb{R}^{d})$, for the solution $u$ of
@@
[…]
@@
where we used that the orthogonal projection $\pi^{\bot}$ commutes with the semigroup. For $m\geq 1$, we also define $\operatorname{|\hskip-1.13809pt|\hskip-1.13809pt|}\cdot\operatorname{|\hskip-1.13809pt|\hskip-1.13809pt|}_{L^{2}(\omega^{\otimes m})}$ as the associated injective cross norm on ${L}^{2}(\omega^{\otimes m})$,
@@ 3. For $P\subset\llbracket m\rrbracket$ and $k,\ell\in P$, we define the operators
    $S_{k,\ell}h^{P}$ $:=$ $-K(x_{k},x_{\ell})\cdot\nabla_{v_{k}}h^{P},$
    $H_{k}h^{P\cup\{*\}}$ $:=$ $-\int_{\mathbb{X}}K(x_{k},x_{*})\cdot\nabla_{v_{k}}h^{P\cup\{*\}}\,\mathrm{d}z_{*}.$
@@ 2.3. Gibbs relaxation
We show that the uniform-in-$N$ exponential relaxation to Gibbs equilibrium in $L^{2}$ can be obtained as a direct estimate on the BBGKY hierarchy (2.9), using hypocoercivity in form of the ergodic estimates of Lemma 2.1.
@@
Now appealing to (2.7) (with constant $1$) and to (2.2) (with $L_{0}$ replaced by $L_{0}^{*}$ and using the equivalence of norms (2.7)), the claim (2.12) follows.
@@ 2.4. Correlation functions
where the symbol $\operatorname{sym}$}’ stands for the symmetrization of coordinates, and we can write more generally, for all $1\leq m\leq N$,
@@ 2.5. BBGKY hierarchy for correlations
Starting from the BBGKY hierarchy (2.9) for marginals of $F^{N}$, and recalling how marginals can be expanded in terms of correlations (2.15) and vice versa (2.17), we can derive a corresponding BBGKY hierarchy of equations for correlations. We refer for instance to [24, Section 4] for the derivation: straightforward but lengthy algebraic manipulations are needed to collect the different factors in this hierarchy.
@@
Given $K\in L^{\infty}(\mathbb{R}^{2d})^{d}$, consider a weak solution $F^{N}\in C(\mathbb{R}_{+};\mathcal{P}\cap{L}^{1}(\mathbb{X}^{N}))$ of the Liouville equation (1.4). The associated correlation functions $\{G^{N,m}\}_{1\leq m\leq N}$ satisfy the following hierarchy of equations: for all $1\leq m\leq N$,
@@ 2.6. Uniform-in-time propagation of chaos and correlation estimates
$(\partial_{t}-L_{0})(F^{N,1}-\mu)\,=\,\kappa H_{1}\Big(\frac{N-1}{N}F^{N,\{1\}}F^{N,\{*\}}-\mu(z_{1})\mu(z_{*})\Big)+\kappa\frac{N-1}{N}H_{1}G^{N,\{1\}\cup\{*\}}.$
@@ 3.1. Relaxation of correlations
First note that the case $m=1$ follows from uniform-in-$N$ Gibbs relaxation in Proposition 1.1(ii). We now consider $2\leq m\leq N$. Starting from the equation for $G^{N,m}$ in Lemma 2.6, taking the difference with the corresponding stationary equation for $C^{N,m}$, and repeating the estimates leading to (2.18), now rather using that for $m\geq 2$ we have
@@ 4.1. Mean-field relaxation
As $A$ is uniformly convex, cf. (1.3), we recall that the probability measure $\propto e^{-A}$ satisfies a Poincaré inequality, see e.g. [33, Section A.19]: in terms of the weight $\omega=e^{A}$, this reads as follows, for all $h\in C^{\infty}_{c}(\mathbb{R}^{d})$,
@@ 4.2. BBGKY hierarchy
$S_{k,\ell}h^{P}$ $:=$ $-{\operatorname{div}}_{x_{k}}(K(x_{k},x_{\ell})h^{P}),$
$H_{k}h^{P\cup\{\ast\}}$ $:=$ $-{\operatorname{div}}_{x_{k}}\Big(\int_{\mathbb{R}^{d}}K(x_{k},x_{*})h^{P\cup\{\ast\}}\,\mathrm{d}x_{*}\Big).$ (4.3)
In this new setting, now using the ergodic estimates in Lemma 4.1 instead of Lemma 2.1, the proof of Proposition 1.1 is immediately adapted. In this way, we slightly extend the result of [34] on uniform-in-time correlation estimates by treating the whole space with confinement, and we obtain a uniform-in-$N$ Gibbs relaxation result in $L^{2}$ that partially improves previous works on the topic.
@@ 5.1. Ergodic estimates for mean-field operators
$L_{M}h\,=\,\left\{\begin{array}[]{lll}\triangle h&:&\text{in Case\penalty 10000\ 1},\\ \triangle h+\kappa\triangle W\ast h&:&\text{in Case\penalty 10000\ 2}.\end{array}\right$.
@@ 5.3. Relaxation of correlations
With the above tools at hand, we now turn to the proof of Theorem 1.6, and we start by adapting Proposition 3.1. Instead of a direct bootstrap argument, we now need to proceed by induction. There is also another difference: uniform-in-$N$ Gibbs relaxation is no longer known to hold a priori (in contrast with Proposition 1.1(ii) in the small-$\kappa$ setting). We thus need to assume that some weak form of Gibbs relaxation holds a priori, which we formulate for convenience in terms of a cross error estimate, cf. (5.2). Note that (5.2) is already known to hold at least for $\theta_{0}=1$ as a consequence of Proposition 5.2 and Corollary 5.3.
@@ 5.4. Cross error estimate for many-particle densities
Acknowledgements
AB acknowledges financial support from the AAP “Accueil” from Université Lyon 1 Claude Bernard, and MD from the European Union (ERC, PASTIS, Grant Agreement n101075879). [^2] Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.